OpenStudy (i_love_my_nieces):

2, 1, 1/2, 1/4, . . . Arithmetic Geometric Both Neither

undeadknight26 (undeadknight26):

Geometric because it goes down by .5 each time...at least how i see it...

OpenStudy (anonymous):

is the difference constant (the same)

OpenStudy (anonymous):

@undeadknight26 please read the code of conduct. do not give answers

OpenStudy (whpalmer4):

If dividing one term by the previous term always gives the same result, it is geometric. If subtracting one term from the previous term always gives the same result, it is arithmetic.

OpenStudy (whpalmer4):

@undeadknight26 The sequence 2, 1.5, 1.0, 0.5, 0 "goes down by .5 each time", also. It isn't geometric, however.

OpenStudy (anonymous):

@whpalmer4 please look a little closer at the sequence provided

OpenStudy (i_love_my_nieces):

Wait I'm confused....?

OpenStudy (whpalmer4):

Lest there be any confusion, when I said "It isn't geometric, however" in my previous post, I was referring only to my example sequence, and not saying anything one way or the other about the one in this problem.

OpenStudy (i_love_my_nieces):

I think it is B, or C.

OpenStudy (whpalmer4):

Test for arithmetic sequence: 2-1 = 1-1/2 = 1/2 - 1/4 = Are they all identical? If so, the sequence is arithmetic Test for geometric sequence: 2/1 = 1/(1/2) = (1/2)/(1/4) = Are they all identical? If so, the sequence is geometric

OpenStudy (i_love_my_nieces):

Yes?

OpenStudy (i_love_my_nieces):

@whpalmer4

OpenStudy (anonymous):

yes to which?

OpenStudy (i_love_my_nieces):

I think it is Geometric.

OpenStudy (i_love_my_nieces):

@whpalmer4 and @pgpilot326

OpenStudy (anonymous):

okay, why?

OpenStudy (i_love_my_nieces):

Idk, am I right?

OpenStudy (anonymous):

idk, if you can't show why then what's the point? you here to get your questions answered or to learn?

OpenStudy (whpalmer4):

Gave you a very simple test. Apply it, and choose. No need to ask us if your answer is correct.

OpenStudy (i_love_my_nieces):

Both.

OpenStudy (anonymous):

lol... fishing?

OpenStudy (i_love_my_nieces):

?